OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x)=5x+1/9x-2

OpenStudy (anonymous):

hi, any ideas?

OpenStudy (anonymous):

Not really no.

OpenStudy (anonymous):

look at the degrees of the polynomials in the numerator and the denominator. what are they?

OpenStudy (anonymous):

what is the polynomial in the numerator?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

and the denominator?

OpenStudy (anonymous):

The same

OpenStudy (anonymous):

good. when that happens, the horizontal asymptote will be the line

OpenStudy (anonymous):

\[y=\frac{\text{leading coefficient of numerator's polynomial} }{ \text{leading coefficient of denominator's polynomial} }\]

OpenStudy (anonymous):

The line?

OpenStudy (anonymous):

Oh so it will be 5/9?

OpenStudy (anonymous):

yep! now that ONLY works if the numerator and denominator have the same degree polynomials

OpenStudy (anonymous):

What happens if they don't?

OpenStudy (anonymous):

then you get different situations... if \[f(x) = \frac{ p(x) }{ q(x) }\] and the degree of p(x) is 1 more than the degree of q(x), you get a slant asymptote and it will be of the form y=ax+b where a \(\ne\) 0. if the degree of q(x) is greater than p(x) then you'll get a horizontal asymptote of y = 0. if f(x) has a different form, say \[f(x)=\frac{ p(x) }{ q(x) }+c\]where c is a non-zero constant and the degree of q(x) is less than p(x), then you get a horizontal asymptote of the form y = c.

OpenStudy (anonymous):

I see. Thanks a lot for the help, I understand it now.

OpenStudy (anonymous):

you're welcome!